OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

Solve the following for x:\[\frac{-3-6}{10-x}=-\frac{3}{2} \]

OpenStudy (anonymous):

so 8

OpenStudy (anonymous):

Verify that x is B.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Plug in 4 for x and then simplify the resulting fraction.

OpenStudy (anonymous):

10-4 eqauls 6

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

\[\frac{-3-6}{10-x} \]\[\frac{-3-6}{10-4}=-\frac{9}{6}=-\frac{3}{2} \]

OpenStudy (anonymous):

4 is the answer isnt it

OpenStudy (anonymous):

it was 4